# DCAD-2000: A Multilingual Dataset across 2000+ Languages with Data Cleaning as Anomaly Detection

### v3

@@ Abstract
The rapid development of multilingual large language models (LLMs) highlights the need for high-quality, diverse, and clean multilingual datasets. In this paper, we introduce DCAD-2000 (Data Cleaning as Anomaly Detection), a large-scale multilingual corpus built using newly extracted Common Crawl data and existing multilingual datasets. DCAD-2000 includes over 2,282 languages, 46.72TB of data, and 8.63 billion documents, spanning 155 high- and medium-resource languages and 159 writing scripts. To overcome the limitations of current data cleaning methods, which rely on manual heuristic thresholds, we propose reframing data cleaning as an anomaly detection task. This dynamic filtering approach significantly enhances data quality by identifying and removing noisy or anomalous content. By fine-tuning LLMs on DCAD-2000, we demonstrate the effectiveness of the data quality improvements, the robustness of the data cleaning pipeline, and enhanced performance for low-resource languages across multiple multilingual benchmarks.
@@ 1 Introduction
Large language models (LLMs) have achieved great progress on a variety of NLP tasks by leveraging vast amounts of training data [1]. However, their performance remains heavily biased towards high-resource languages [2, 3]. To improve the multilingual capabilities of LLMs, a common strategy is to incorporate large amounts of non-English data, either by continue pretraining [4] or by instruction tuning in multilingual settings [5]. Therefore, constructing large-scale, high-quality multilingual datasets is crucial for enhancing the multilingual performance of LLMs.
@@
mC4 [6] $C$-MAIN-2020-34 101 0 43 52 6
OSCAR 23.01 [7] $C$-MAIN-2022-49 153 6 42 25 80
Glot500 [8] $C$-MAIN-2020-34 511 0 108 79 324
CulturaX [9] $C$-MAIN-2022-49 167 11 47 27 82
Madlad-400 [10] $C$-MAIN-2022-33 419 7 46 39 327
MaLA [11] $C$-MAIN-2022-49 939 1 125 78 735
Glotcc [12] $C$-MAIN-2023-50 1331 0 10 52 1269
HPLT-v1.2 [13] $C$-MAIN-2022-40 191 12 53 38 88
Fineweb-2 [14] $C$-MAIN-2024-18 1915 10 62 49 1794
@@
Recent efforts to construct multilingual datasets have built several notable resources, including CulturaX [9], HPLT [13], Madlad-400 [10], MaLA [15], and Glotcc [12], which cover 167, 191, 419, 939, and 1,331 languages, respectively. While these datasets have made significant contributions, they exhibit three major limitations, as summarized in Table 1: (1) Outdated data sources: These datasets primarily rely on older versions of Common Crawl [^1] https://commoncrawl.org, which results in outdated knowledge and an elevated risk of hallucination [16]. (2) Limited coverage of high- and medium-resource languages [^2] We follow the criteria from Flores-101 [17] to categorize languages: High: >100M>100M; Medium: $(1M,100M)$; Low: $(100K,1M)$; Very Low: <100K<100K.: For instance, Fineweb-2 [14], which supports 1,915 languages, includes data from only 10 high-resource languages and 62 medium-resource languages. (3) Insufficient data cleaning: Despite being cleaned, recent studies [18, 19] indicate that these datasets still contain a significant amount of noise, which makes them difficult to directly employ in training multilingual LLMs. For instance, Sailor [18] founds that 31.11% of the data in Madlad-400 could still be further removed by advanced data cleaning.
@@
Traditional data cleaning workflows [20] often rely on document-level features (e.g., language identification; 21) and fixed thresholds to filter out low-quality data. However, this approach struggles with cross-lingual consistency due to feature distribution differences. For example [^1] Please refer to Appendix A for more details., the variation in average word counts across languages, datasets, and shards demonstrates the limitations of heuristic threshold-based data cleaning methods. Notably, while Fineweb-2 fine-tunes thresholds for more than 1,000 languages, this process is computationally intensive and time-consuming.
@@
To address these challenges, we introduce DCAD-2000, a new large-scale, high-quality multilingual dataset that can be directly applied to LLM training. DCAD-2000 covers 2282 languages (155 high/medium languages), incorporating the latest Common Crawl data (November 2024; CC-MAIN-2024-46) and existing multilingual datasets. Additionally, we propose a novel language-agnostic data cleaning approach that treats data cleaning as an anomaly detection [22] problem, distinguishing it from traditional threshold-based methods [14, 23]. Our approach extracts eight statistical features from each document to evaluate quality, including number of words, character/word repetition ratio, special character/word ratio, stopword ratio, flagged words ratio, language identification score and perplexity score. Anomaly detection algorithms dynamically identify and remove outliers by recognizing deviations from typical document quality metrics.
@@
We conduct a comprehensive analysis of DCAD-2000 with respect to document distribution, linguistic and geographical characteristics, writing scripts, and resource classification (Section 4). By fine-tuning LLMs on DCAD-2000, we validate the effectiveness of its data quality and data cleaning pipeline. Furthermore, we demonstrate the superiority of DCAD-2000 across various language categories (high, medium, low and very low) in multiple multilingual benchmarks, including SIB-200 [24], Glot500 [8] and FLORES-200 [25] (Section 5).
In summary, we make the following contributions: (1) We propose a novel data cleaning framework that frames the task as anomaly detection, offering a language-agnostic and adaptive solution without manual threshold tuning. (2) We release DCAD-2000, a comprehensive multilingual dataset covering over 2,282 languages, containing 8.63B of documents, 46.72TB of disk size and 159 writing scripts with metadata annotations. (3) Extensive evaluation across multiple multilingual benchmarks demonstrates the effectiveness of both the data quality and the data cleaning pipeline.
@@ 2 Related Work
Enhancing the multilingual capabilities of LLMs often involves continuing pretraining on large-scale multilingual datasets [11, 15]. These datasets can be broadly categorized into curated corpora, domain-specific corpora, and web-crawled corpora. (I) Curated Corpora. Curated datasets are carefully gathered by experts from high-quality sources such as books [23], academic publications [26], and encyclopedia entries [27, 28, 29]. (II) Domain-Specific Corpora. In addition to general-domain data, fine-tuning LLMs on domain-specific datasets is crucial for improving performance in specialized domains like finance [30], healthcare [31], legal [32], and education [33, 34]. (III) Web-Crawled Corpora. Web-crawled datasets, particularly those derived from Common Crawl, provide large-scale multilingual coverage by leveraging an open repository of over 250 billion web pages. These datasets include mC4 [6], CC-100 [35], OSCAR [7], Glotcc [12], Fineweb [36], and Fineweb-2 [14]. While curated and domain-specific corpora offer high-quality content with limited language coverage, web-crawled corpora provide broader multilingual coverage but often suffer from noise and lower data quality [18, 19].
@@
Data cleaning is an essential step in preparing high-quality datasets for training robust LLMs. It involves filtering noisy, irrelevant, or harmful content and can be broadly classified into model-based and heuristic-based approaches [37]. (I) Model-Based Methods. Model-based approaches employ classifiers or LLMs to distinguish between high-quality and low-quality data. For instance, content safety models [38] filter out explicit or gambling-related content, while quality classifiers remove low-relevance text [39]. LLM-based methods focus on generating prompts for cleaning [40] or integrating error detection and correction into the pipeline [41, 42]. (II) Heuristic-Based Methods. Heuristic approaches apply predefined rules to filter content at both document and sentence levels. At the document level, strategies include filtering by language identification scores or scoring documents with language models [9, 23]. At the sentence level, rules are applied to remove incomplete or irrelevant content, such as HTML tags or excessively short sentences [6, 7]. While model-based methods offer high precision but face scalability challenges, heuristic-based methods are more efficient yet less adaptable to diverse multilingual data.
@@ 3.1 Data Collection
To ensure comprehensive and robust multilingual data representation, DCAD-2000 integrates data from four main sources: MaLA, Fineweb, Fineweb-2, and newly extracted Common Crawl data. Each source is selected based on its unique contribution to multilingual coverage, data quality, and freshness, with careful consideration to complementarity to minimize redundancy. Specifically, MaLA and Fineweb-2 are prioritized due to their broad language coverage and high-quality curation, which complements other widely used datasets like mC4 [6] and OSCAR [7].
MaLA Corpus [11]. The MaLA corpus covers 939 languages, aggregating data from diverse sources including Bloom [43], CC100 [35], Glot500 [8], among others. Deduplication is performed using MinHashLSH [44], which is particularly effective in removing near-duplicate entries that often arise from common web sources. Language codes are based on ISO 639-3222https://en.wikipedia.org/wiki/ISO_639-3 standards, and language-specific scripts are supported by GlotScript [^3] https://github.com/cisnlp/GlotScript.
Fineweb Corpus [36]. Fineweb is a high-quality English web dataset extracted from Common Crawl, consisting of over 15 trillion tokens and updated monthly. Data cleaning and deduplication are performed using the Datatrove library. [^4] https://github.com/huggingface/datatrove For DCAD-2000, we incorporate data from the November 2024 release (CC-MAIN-2024-46) to ensure freshness and up-to-date relevance of the data.
Fineweb-2 Corpus [14]. Fineweb-2 expands Fineweb to include multilingual data, covering 1,915 languages. It processes 96 Common Crawl dumps from 2013 (CC-MAIN-2013-20) to April 2024 (CC-MAIN-2024-20). The deduplication process within Fineweb-2 is similarly handled using the Datatrove library, ensuring the exclusion of redundant entries and maintaining high-quality multilingual coverage.
@@ 3.2.1 Feature Extraction
Inspired by Roots [23] and CulturaX [9], we extract eight statistical features from each document to evaluate text quality. Each feature is selected for its ability to capture important characteristics of the text, contributing to robust anomaly detection. Let $t$ represent a document; the extracted features are: (1) Number of Words, $n_{w}(t)$: Total token count after tokenization. (2) Character Repetition Ratio, $r_{c}(t)$: Fraction of repeated character sequences, highlighting noise or encoding errors. (3) Word Repetition Ratio, $r_{w}(t)$: Proportion of repeated words, indicating redundancy or unnatural text. (4) Special Characters Ratio, $r_{s}(t)$: Fraction of special characters based on language-specific lists from ROOTs Corpus [23]. These lists include emojis, whitespace types, numbers, and punctuation. (5) Stopwords Ratio, $r_{\text{stop}}(t)$: Proportion of stopwords using Fineweb-2’s language-specific stopword lists. (6) Flagged Words Ratio, $r_{\text{flag}}(t)$: Fraction of toxic or profane words from Toxicity-200 [25] or public repositories [^6] https://github.com/thisandagain/washyourmouthoutwithsoap. (7) Language Identification (LID) Score, $s_{\text{lid}}(t)$: Confidence score from GlotLID [21], supporting over 2,000 languages. (8) Perplexity Score, $s_{\text{ppl}}(t)$: Using KenLM [45], we train a language model for each language on multilingual Wikipedia data (November 2023) [^7] For languages not covered by the trained models, a default perplexity score of 500 is assigned..
@@ 3.2.2 Anomaly Detection
where $\tau\in\mathbb{R}$ is a hyperparameter determined empirically or through cross-validation. [^9] We use the default setting of the specific anomaly detection algorithm.
@@ 3.3 Computational Resources
The construction of the DCAD-2000 dataset leveraged Ksyun servers [^10] https://k8s.ksyun.com to process and clean the multilingual data efficiently. Each server instance is equipped with 32 CPU cores, 128GB of memory, and 100GB of disk storage, which is utilized for intermediate data handling and memory-intensive operations such as anomaly detection. The workload is managed using container orchestration tools, Kubernetes [^11] https://kubernetes.io, with up to 100 parallel tasks running per job to ensure scalability.
@@ 4 Dataset Analysis
Language Resource Classification. Following the classification approach proposed by Flores [17], we categorize languages in DCAD-2000 into four groups based on corpus size: high-resource, medium-resource, low-resource, and extremely low-resource. Table 1 shows the distribution across these categories. The dataset includes 155 high- and medium-resource languages, while low-resource languages make up a significant portion, which reflects DCAD-2000’s commitment to supporting underrepresented languages. Notably, DCAD-2000 surpasses other corpora in its balance between high-resource and low-resource languages, which can have a significant impact on multilingual model training. The distribution of languages across categories ensures that the dataset is well-suited for developing models that perform effectively across diverse language resources.
Impact of Data Cleaning. We summarize the document count, token count, and disk size of the high/medium/low resource languages in DCAD-2000 before and after the data cleaning process. Complete details are provided in Appendix C. The cleaning process results in the removal of a substantial amount of noisy data, even from datasets like MaLA, Fineweb, and Fineweb-2, which had already been subject some cleaning. This aligns with the findings from [18, 19]. For example, in the MaLA dataset, 8.05 million documents are removed for the hbs_Latn language, which suggests the necessity of rigorous data cleaning to enhance dataset quality. Overall, the cleaning process removed approximately 7.69% of the documents across all languages, significantly improving the quality of the dataset by reducing noise and increasing relevance for model training (Section 5).

### v4

@@ Abstract
The rapid development of multilingual large language models (LLMs) highlights the need for high-quality, diverse, and well-curated multilingual datasets. In this paper, we introduce DCAD-2000 (Data Cleaning as Anomaly Detection), a large-scale multilingual corpus constructed from newly extracted Common Crawl data and existing multilingual sources. DCAD-2000 covers 2,282 languages, 46.72TB of text, and 8.63 billion documents, spanning 155 high- and medium-resource languages and 159 writing scripts. To overcome the limitations of existing data cleaning approaches, which rely on manually designed heuristic thresholds, we reframe data cleaning as an anomaly detection problem. This dynamic filtering paradigm substantially improves data quality by automatically identifying and removing noisy or anomalous content. By fine-tuning LLMs on DCAD-2000, we demonstrate notable improvements in data quality, robustness of the cleaning pipeline, and downstream performance, particularly for low-resource languages across multiple multilingual benchmarks.
@@ 1 Introduction
Large language models (LLMs) have achieved great progress on a variety of NLP tasks by leveraging vast amounts of training data minaee2024large . However, their performance remains heavily biased towards high-resource languages zhu2024multilingual ; huang2024survey . To improve the multilingual capabilities of LLMs, a common strategy is to incorporate large amounts of non-English data, either by continue pretraining lai-etal-2024-llms or by instruction tuning in multilingual settings ustun-etal-2024-aya . Therefore, constructing large-scale, high-quality multilingual datasets is crucial for enhancing the multilingual performance of LLMs.
@@
mC4 raffel2020exploring $C$-MAIN-2020-34 101 0 43 52 6
OSCAR 23.01 abadji2022towards $C$-MAIN-2022-49 153 6 42 25 80
Glot500 imanigooghari-etal-2023-glot500 $C$-MAIN-2020-34 511 0 108 79 324
CulturaX nguyen-etal-2024-culturax $C$-MAIN-2022-49 167 11 47 27 82
Madlad-400 kudugunta2024madlad $C$-MAIN-2022-33 419 7 46 39 327
MaLA ji2024emma $C$-MAIN-2022-49 939 1 125 78 735
Glotcc kargaran2024glotcc $C$-MAIN-2023-50 1331 0 10 52 1269
HPLT-v1.2 de-gibert-etal-2024-new $C$-MAIN-2022-40 191 12 53 38 88
Fineweb-2 penedo2025fineweb2 $C$-MAIN-2024-18 1915 10 62 49 1794
@@
Recent efforts have introduced several large multilingual corpora, including CulturaX nguyen-etal-2024-culturax , HPLT de-gibert-etal-2024-new , Madlad-400 kudugunta2024madlad , MaLA lin2024mala , and Glotcc kargaran2024glotcc , which cover 167, 191, 419, 939, and 1,331 languages, respectively. While these datasets have made significant contributions, they exhibit three major limitations, as summarized in Table 1: (1) Outdated data sources: These datasets primarily rely on older Common Crawl snapshots [^1] https://commoncrawl.org, which results in outdated knowledge and an elevated risk of hallucination huang2023survey . (2) Limited coverage of high- and medium-resource languages [^2] We follow the criteria from Flores-101 goyal-etal-2022-flores to categorize languages: High: >100M>100M; Medium: $(1M,100M)$; Low: $(100K,1M)$; Very Low: <100K<100K.: For instance, Fineweb-2 penedo2025fineweb2 , despite supporting 1,915 languages, contains data from only 10 high-resource and 62 medium-resource languages. (3) Insufficient data cleaning: Despite being cleaned, recent studies dou-etal-2024-sailor ; zhang-etal-2024-mc2 indicate that these datasets still contain a significant amount of noise, which makes them difficult to directly employ in training multilingual LLMs. For example, Sailor dou-etal-2024-sailor reports that 31.11% of Madlad-400 data could still be removed using more advanced cleaning.
@@
Traditional data cleaning workflows albalak2024survey often rely on document-level heuristics (e.g., language identification; kargaran-etal-2023-glotlid, ) and fixed thresholds to filter low-quality samples. However, these heuristic thresholds often fail to generalize across languages due to distributional differences in features such as word count, repetition ratios, and perplexity [^1] Please refer to Appendix A for more details.. Notably, while Fineweb-2 fine-tunes thresholds for more than 1,000 languages, this process is computationally intensive and time-consuming.
@@
To address these challenges, we introduce DCAD-2000, a new large-scale, high-quality multilingual dataset that can be directly applied to LLM training. DCAD-2000 covers 2282 languages (155 high/medium languages), incorporating the latest Common Crawl data (November 2024; CC-MAIN-2024-46) and existing multilingual datasets. Additionally, we propose a novel language-agnostic data cleaning approach that treats data cleaning as an anomaly detection su2024large problem, distinguishing it from traditional threshold-based methods penedo2025fineweb2 ; laurenccon2022bigscience . Our approach extracts eight statistical features, including number of words, character/word repetition ratio, special character/word ratio, stopword ratio, flagged words ratio, language identification score and perplexity score. Anomaly detection algorithms dynamically identify and remove outliers by recognizing deviations from typical document quality metrics.
@@
We conduct a comprehensive analysis of DCAD-2000 with respect to document distribution, linguistic and geographical characteristics, writing scripts, and resource classification (Section 4). By fine-tuning LLMs on DCAD-2000, we validate the effectiveness of its data quality and data cleaning pipeline. Furthermore, we demonstrate the superiority of DCAD-2000 across various language categories (high, medium, low and very low) in multiple multilingual benchmarks, including SIB-200 adelani-etal-2024-sib , Glot500 imanigooghari-etal-2023-glot500 and FLORES-200 costa2022no (Section 5).
In summary, we make the following contributions:
  • We propose a novel data cleaning framework that frames the task as anomaly detection, offering a language-agnostic and adaptive solution without manual threshold tuning.
  • We release DCAD-2000, a comprehensive multilingual dataset covering over 2,282 languages, containing 8.63B of documents, 46.72TB of disk size and 159 writing scripts with metadata annotations.
  • Extensive evaluation across multiple multilingual benchmarks demonstrates the effectiveness of both the data quality and the data cleaning pipeline.
@@ 2 Related Work
Enhancing the multilingual capabilities of LLMs often involves continuing pretraining on large-scale multilingual datasets ji2024emma ; lin2024mala . These datasets can be broadly categorized into curated corpora, domain-specific corpora, and web-crawled corpora. (I) Curated Corpora. Curated datasets are carefully gathered by experts from high-quality sources such as books laurenccon2022bigscience , academic publications clement2019use , and encyclopedia entries brown2011wikipedia ; lehmann2015dbpedia ; kuo2024wikibench . (II) Domain-Specific Corpora. In addition to general-domain data, fine-tuning LLMs on domain-specific datasets is crucial for improving performance in specialized domains like finance zhang2023xuanyuan , healthcare wu2024pmc , legal colombo2024saullm , and education xu2024large ; lozhkov2024fineweb-edu . (III) Web-Crawled Corpora. Web-crawled datasets, particularly those derived from Common Crawl, provide large-scale multilingual coverage by leveraging an open repository of over 250 billion web pages. These datasets include mC4 raffel2020exploring , CC-100 conneau-etal-2020-unsupervised , OSCAR abadji2022towards , Glotcc kargaran2024glotcc , Fineweb penedo2024fineweb , and Fineweb-2 penedo2025fineweb2 . While curated and domain-specific corpora offer high-quality content with limited language coverage, web-crawled corpora provide broader multilingual coverage but often suffer from noise and lower data quality dou-etal-2024-sailor ; zhang-etal-2024-mc2 .
@@
Data cleaning is an essential step in preparing high-quality datasets for training robust LLMs. It involves filtering noisy, irrelevant, or harmful content and can be broadly classified into model-based and heuristic-based approaches liu2024datasets . (I) Model-Based Methods. Model-based approaches employ classifiers or LLMs to distinguish between high-quality and low-quality data. For instance, content safety models li-etal-2024-salad filter out explicit or gambling-related content, while quality classifiers remove low-relevance text jiang2024more . LLM-based methods focus on generating prompts for cleaning narayan2022can or integrating error detection and correction into the pipeline chen2023seed ; ni2024iterclean . (II) Heuristic-Based Methods. Heuristic approaches apply predefined rules to filter content at both document and sentence levels. At the document level, strategies include filtering by language identification scores or scoring documents with language models nguyen-etal-2024-culturax ; laurenccon2022bigscience . At the sentence level, rules are applied to remove incomplete or irrelevant content, such as HTML tags or excessively short sentences raffel2020exploring ; abadji2022towards . While model-based methods offer high precision but face scalability challenges, heuristic-based methods are more efficient yet less adaptable to diverse multilingual data.
@@ 3.1 Data Collection
To ensure comprehensive and robust multilingual data representation, DCAD-2000 integrates data from four main sources: MaLA, Fineweb, Fineweb-2, and newly extracted Common Crawl data. Each source is selected based on its unique contribution to multilingual coverage, data quality, and freshness, with careful consideration to complementarity to minimize redundancy. Specifically, MaLA and Fineweb-2 are prioritized due to their broad language coverage and high-quality curation, which complements other widely used datasets like mC4 raffel2020exploring and OSCAR abadji2022towards .
MaLA Corpus ji2024emma . The MaLA corpus covers 939 languages, aggregating data from diverse sources including Bloom leong-etal-2022-bloom , CC100 conneau-etal-2020-unsupervised , Glot500 imanigooghari-etal-2023-glot500 , among others. Deduplication is performed using MinHashLSH broder1998min , which is particularly effective in removing near-duplicate entries that often arise from common web sources. Language codes are based on ISO 639-3222https://en.wikipedia.org/wiki/ISO_639-3 standards, and language-specific scripts are supported by GlotScript [^3] https://github.com/cisnlp/GlotScript.
Fineweb Corpus penedo2024fineweb . Fineweb is a high-quality English web dataset extracted from Common Crawl, consisting of over 15 trillion tokens and updated monthly. Data cleaning and deduplication are performed using the Datatrove library. [^4] https://github.com/huggingface/datatrove For DCAD-2000, we incorporate data from the November 2024 release (CC-MAIN-2024-46) to ensure freshness and up-to-date relevance of the data.
Fineweb-2 Corpus penedo2025fineweb2 . Fineweb-2 expands Fineweb to include multilingual data, covering 1,915 languages. It processes 96 Common Crawl dumps from 2013 (CC-MAIN-2013-20) to April 2024 (CC-MAIN-2024-20). The deduplication process within Fineweb-2 is similarly handled using the Datatrove library, ensuring the exclusion of redundant entries and maintaining high-quality multilingual coverage.
@@ 3.2.1 Feature Extraction
Inspired by Roots laurenccon2022bigscience and CulturaX nguyen-etal-2024-culturax , we extract eight statistical features from each document to evaluate text quality. Each feature is selected for its ability to capture important characteristics of the text, contributing to robust anomaly detection. Let $t$ represent a document; the extracted features are:
  • Number of Words, $n_{w}(t)$: Total number of tokens after language-specific tokenization, providing a coarse measure of document length and helping identify extremely short or excessively long outliers.
  • Character Repetition Ratio, $r_{c}(t)$: Fraction of repeated character sequences (e.g., “aaaaa” or “!!!!!”), which often signal encoding artifacts, copy-paste errors, or spam-like content.
  • Word Repetition Ratio, $r_{w}(t)$: Proportion of repeated lexical items, useful for detecting low-information documents that exhibit looping or template-like patterns.
  • Special Characters Ratio, $r_{s}(t)$: Fraction of characters belonging to special symbol categories. We employ the curated language-specific symbol lists provided in the ROOTs Corpus laurenccon2022bigscience , covering punctuation, numeric symbols, whitespace variants, and emojis. A high $r_{s}(t)$ may indicate adversarial inputs or unstructured noise.
  • Stopwords Ratio, $r_{\text{stop}}(t)$: Ratio of stopwords derived from Fineweb-2’s multilingual stopword lexicons. This metric captures the functional-to-content word balance, offering a lightweight approximation of linguistic naturalness.
  • Flagged Words Ratio, $r_{\text{flag}}(t)$: Fraction of tokens that appear in curated lists of toxic or profane vocabulary such as Toxicity-200 costa2022no and community-maintained sources [^6] https://github.com/thisandagain/washyourmouthoutwithsoap. This feature enables early detection of harmful or sensitive content.
  • Language Identification (LID) Score, $s_{\text{lid}}(t)$: Confidence score produced by GlotLID kargaran-etal-2023-glotlid , a language identifier supporting over 2,000 languages. Lower scores may indicate code-switching, mislabeling, or mixed-script anomalies.
  • Perplexity Score, $s_{\text{ppl}}(t)$: We compute a language model perplexity score using KenLM heafield-2011-kenlm models trained per language on the November 2023 snapshot of multilingual Wikipedia [^7] KenLM models are only trained for languages with sufficient clean Wikipedia data (minimum 10,000 high-quality sentences). For other languages, we assign a default perplexity score of 500.. This feature provides a lightweight proxy for linguistic fluency.
@@ 3.2.2 Anomaly Detection
where $\tau\in\mathbb{R}$ is a hyperparameter determined empirically or through cross-validation. [^9] We use the default settings of the specific anomaly detection algorithm in Scikit-learn, applying these settings globally rather than individually for each feature or language.
@@ 3.3 Computational Resources
The construction of the DCAD-2000 dataset leveraged Ksyun servers [^10] https://www.ksyun.com to process and clean the multilingual data efficiently. Each server instance is equipped with 32 CPU cores, 128GB of memory, and 100GB of disk storage, which is utilized for intermediate data handling and memory-intensive operations such as anomaly detection. The workload is managed using container orchestration tools, Kubernetes [^11] https://kubernetes.io, with up to 100 parallel tasks running per job to ensure scalability.
@@ 4 Dataset Analysis
Language Resource Classification. Following the classification approach proposed by Flores goyal-etal-2022-flores , we categorize languages in DCAD-2000 into four groups based on corpus size: high-resource, medium-resource, low-resource, and extremely low-resource. Table 1 shows the distribution across these categories. The dataset includes 155 high- and medium-resource languages, while low-resource languages make up a significant portion, which reflects DCAD-2000’s commitment to supporting underrepresented languages. Notably, DCAD-2000 surpasses other corpora in its balance between high-resource and low-resource languages, which can have a significant impact on multilingual model training. The distribution of languages across categories ensures that the dataset is well-suited for developing models that perform effectively across diverse language resources.
Impact of Data Cleaning. We summarize the document count, token count, and disk size of the high/medium/low resource languages in DCAD-2000 before and after the data cleaning process. Complete details are provided in Appendix C. The cleaning process results in the removal of a substantial amount of noisy data, even from datasets like MaLA, Fineweb, and Fineweb-2, which had already been subject some cleaning. This aligns with the findings from dou-etal-2024-sailor ; zhang-etal-2024-mc2 . For example, in the MaLA dataset, 8.05 million documents are removed for the hbs_Latn language, which suggests the necessity of rigorous data cleaning to enhance dataset quality. Overall, the cleaning process removed approximately 7.69% of the documents across all languages, significantly improving the quality of the dataset by reducing noise and increasing relevance for model training (Section 5).